The speed of a stream is 4mph. If a boat travels 80 miles downstream in the same time that it takes to travel 40 miles upstream, what is the speed of the boat in still water?
Answer by checkley77(12569)

D=RT
80=(D+4)T
40=(D-4)T
SEEING AS THE TIMES ARE THE SAME WE HAVE:
80/(D+4)=40/(D-4)
80(D-4)=40(D+4)
80D-320=40D+160
80D-40D=160+320
40D=480
D=480/40
D=12 MPH FOR THE CURRENT.
PROOF;
80/(12+4)=40/(12-4)
80/16=40/8
16*40=80*8
640=640
----------------------------------------------------------
Hi Thank you for your time. One that I simply cannot get is: An airplane
travels 600 miles against the wind in 4 hrs. and makes the return trip with the
same wind in 2 hrs. Find the rate of the wind. The answer is either: 75, 150,
225, or 300mph. I can't seem to get the formula and it is driving me crazy!
Can you help?
D=R*T
600=R*4
R=600/4=150 MPH AGAINST THE WIND.
600=R*2
R=600/2=300 MPH WITH THE WIND.
(300-150)/2=150/2=75 MPH IS THE SPEED OF THE WIND.
-------------------------------------------------